Suppose you did a chi-square procedure with 50 subjects where there were four categories. What would be the degrees of freedom?
Answer #1

If there is K subjects and if each subjects has r categories then the df of chi square distribution is = (k-1)(r-1)

Here k = 50 , r = 4

So df = 49×3 = 147